40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15554     Entry time: Thu Sep 3 00:00:57 2020
Author: gautam 
Type: Update 
Category: BHD 
Subject: New patch cable installed 
  • 10m PM1064 cable was installed. I tried a double shielding approach (photos tmrw here), but I suspect the real weak point is where the fiber is plugged into the collimator - it's hard to imagine we can stabilize this sort of arrangement to better than 100um absolute length over long periods of time, I'd think thermal/mechanical strains in the fiber will modulate the length by ~mm (?). Anyways, let's see what the heterodyne measurement tells us.
  • This work required (i) realignment at the input coupler and (ii) change of position of mode matching lenses in the LO path on the AS table to see any interference with the IFO beam. This indicates something was seriously wrong with the old patch cable, as the collimator should set the mode. The MFD of the two fibers may have been different, but I don't know if that alone can account for it.
  • As of now, I have fringes between the ITM single bounce and the LO, but the fringe pk-pk is only 10% of the theoretical pk-pk based on DC values of the LO and AS beams. So the mode matching can be improved significantly (I preivously had ~60%).

To be continued tomorrow. I think it's a good idea to let the newly installed fiber relax into some sort of stable configuration overnight.

ELOG V3.1.3-